OpenStudy (anonymous):

MEDAL AND FOLLOW For the function f(x) = x+7 what is the ordered pair for the point on the graph when x = 2b

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Hi, Austin! If f(x)=x+7, and you want to know the value of this function when x=2, you'd just substitute 2 for x. Would you do that now? Then we'll move on to doing the rest of the problem.

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

I i know how to do the problem but not with a graph

OpenStudy (anonymous):

@mathmale is it a?

OpenStudy (anonymous):

A : 2b ,2b +7

OpenStudy (anonymous):

?

OpenStudy (mathmale):

Austin, you didn't type out your answer choices, so I can only guess what your answer 'A' looks like. At any rate, I'm always more interested in knowing HOW you got your answer than I am in the answer itself. If you have the function f(x)=x+7 and (arbitrarily) let x=a, then f(a)=a+7. this says, "the value of the function f(x), when x=a, is f(a)=a+7. Now: would you please find f(b)? Type out your answer. Next, would you please find f(2b)?

OpenStudy (mathmale):

The topic here, in case you're wondering, is "evaluating functions."